OpenStudy (anonymous):

tan3x=-1 , solve for all solutions

OpenStudy (mertsj):

\[3x=\frac{\pi}{4}, \frac{5\pi}{4}, \frac{9\pi}{4}, \frac{13\pi}{4}, \frac{17\pi}{4}, \frac{21\pi}{4}\]

OpenStudy (mertsj):

\[x=\frac{\pi}{12}, \frac{5\pi}{12}, \frac{9\pi}{12}, \frac{13\pi}{12}, \frac{17\pi}{12}, \frac{21\pi}{12}+2\pi n\]

OpenStudy (anonymous):

would the answers be the same if the q was solve for cot(3x)=-1

OpenStudy (mertsj):

Oh boy. I didn't see the negative sign. The answers for 3x should be in quadrants 2 and 4 instead of 1 and 3. So replace the numerators with these numbers:

OpenStudy (mertsj):

\[3\pi, 7\pi,11\pi, 15\pi, 19\pi, 23\pi\]

OpenStudy (mertsj):

And yes, since the cot is the reciprocal of the tangent, and since the reciprocal of -1 is -1, the answers would be the same.

OpenStudy (anonymous):

can u help me with another one

OpenStudy (mertsj):

Post it.

OpenStudy (anonymous):

kay here it is csc^2(2x)-csc(2x)-2=0

OpenStudy (mertsj):

\[\csc ^2(2x)-\csc (2x)-2=0\]

OpenStudy (mertsj):

\[(\csc (2x)-2)(\csc (2x)+1)=0\]

OpenStudy (mertsj):

\[\csc (2x)=2 , \csc (2x)=-1\]

OpenStudy (mertsj):

\[2x=\frac{\pi}{6}, \frac{5\pi}{6}, \frac{13\pi}{6}, \frac{17\pi}{6}\]

OpenStudy (mertsj):

\[x=\frac{\pi}{12}, \frac{5\pi}{12}, \frac{13\pi}{12}, \frac{17\pi}{12} + 2\pi n\]

OpenStudy (mertsj):

\[\csc (2x)=-1\]

OpenStudy (anonymous):

oh and can u help me with a true and false question for this equeation: 3tan^2(2x)-1=0

OpenStudy (anonymous):

I already got the question above lets move on

OpenStudy (anonymous):

This equation has 8 solutions on the domain . is one of the solutions of this equation. The general solution of this equation is . is one of the solutions of this equation which option is not true

OpenStudy (mertsj):

\[2x=\frac{3\pi}{2}, \frac{7\pi }{2}, x=\frac{3\pi}{4}, \frac{7\pi}{4} + 2\pi n\]

OpenStudy (anonymous):

its missing some info

OpenStudy (mertsj):

So how am I supposed to answer it?

OpenStudy (anonymous):

so for the first option it was supposed to say in the domain of 0 to 2pi

OpenStudy (anonymous):

the second option was supposed to say 13pi/12 is one of the solutions

OpenStudy (mertsj):

8 solutions.

OpenStudy (anonymous):

so the eight solutions one is the false one

OpenStudy (anonymous):

one last question for you

OpenStudy (mertsj):

I cannot tell which one is the false one because the answers are not posted.